A shell is fired from the ground with an initial speed of 1.70 × 103 m/s (approximately five
times the speed of sound) at an initial angle of 52.0 degrees to the horizontal.
The acceleration of gravity is 9.81 m/s^2
a) Neglecting air resistance, find the shell’s horizontal range.
1
Expert's answer
2018-09-17T15:37:08-0400
x=v_0 cos⁡α t
v_0 sin⁡α t-(gt^2)/2=0
x=(v_0 cos⁡α 2v_0 sin⁡α)/g=(2v_0^2 sin⁡2α)/g=(2·〖(1.7·〖10〗^3)〗^2·sin⁡〖104 ̊〗)/9.81=571519 m=571.52 km